1. Automobiles and trucks pollute the air with NO. At 2000.0°C, Kc for the reaction N2(g)+O2(g)⇌2NO(g) At 2000.0°C Kc = 4.100 × 10–4 and ΔH° = 180.6 kJ. What is the value of Kc at 83.00°C?
1
Expert's answer
2020-04-10T13:35:17-0400

Use Van't Hoff equation

"ln\frac{K_2}{K_1}=\frac{-\Delta H}{R}(\frac{1}{T_2}-\frac{1}{T_1})"


"T_1(K)= 2000.0 + 273= 2273 K"

"T_2(K) = 83.00+273=356.K"

"\Delta H = 180.6 kJ = 180600J"

"K_1 = 4.100\times10^{-4}"

"K_2 - ?"


"ln\frac{K_2}{4.100\times 10^{-4}}= \frac {-180600}{8.314}(\frac{1}{356}-\frac{1}{2273})"


"K_2 = 1.836\times10^{-26}"